HETAV PATEL

Grade 11,

A square and a rectangle have the same perimeter. Calculate the area of the rectangle if the side of the square is 60 cm and the length of the rectangle is 80 cm.

A square and a rectangle have the same perimeter. Calculate the area of the rectangle if the side of the square is 60 cm and the length of the rectangle is 80 cm.

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Perimeter of square formula = 4× side of the square

Hence, P (square) = 4 × 60 = 240 cm

Perimeter of rectangle formula = 2× (Length + Breadth)

Hence, P (rectangle) = 2 (80 + Breadth)

= 160 + 2 × Breadth

According to the given question,
160 + 2× Breadth = 240 cm
2× Breadth = 240 – 160
Breadth = 80/2
The breadth of the rectangle = 40 cm

Now, the area of rectangle = Length × Breadth = 80 × 40 = 3200 cm^2

Thanks